Two angles of triangle measures 35 and 90 degrees. the sum of triangles of any triangle equal 180 degrees. Uses this information to write an equation to solve for the third angle of the triangle and solve it.

Answer by TimothyLamb(4379)

---
35 + 90 + x = 180
x = 180 - (35 + 90)
x = 180 - 125
x = 55
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php